Twenty-two years ago, Mark was twice as old as Bren and eight times as old as Ron. The same year, the sum of their ages was 26. How old is Bren now?
Answer by ikleyn(52818)

.

22 years ago:


    Ron  =  x

    Mark = 8x

    Bren = 4x


    The sum of ages 22 years ago  was


        x + 8x + 4x = 26

        13x         = 26

          x         = 26/13 = 2.


22 years ago, Bren was 4x = 4*2 = 8 years old.


Now, Bren is 8 + 22 = 30 years old.     ANSWER

Solved.